OpenStudy (anonymous):

a man is 65 5/9 tall his daughter 59 5/12 taller how much taller is the man?

OpenStudy (timaashorty):

\[ 65 \frac{ 5 }{ 9 } - 65 \frac{ 5 }{ 12} = 6 \frac{ 5 }{ 36 }\]

OpenStudy (anonymous):

omg thank uuuuuuuu

OpenStudy (timaashorty):

You're welcome

OpenStudy (timaashorty):

MEdal please ? [: